\begin{thebibliography}{HD82}

\normalsize
\baselineskip=15pt

%%%%%%%%%%%%%
\bibitem[Abr59]{a59} L. Abramov, On the entropy of a flow, \emph{Dokl. Akad. Nauk SSSR} \textbf{128} (1959), 873-875.

\bibitem[Bow73]{b73} R. Bowen,  Topological entropy for noncompact sets,  \emph {Trans. Amer. Math. Soc.} \textbf{184} (1973), 125-136.

\bibitem[BR75]{br75} R. Bowen and  D. Ruelle, The ergodic theory of Axiom A flows, \emph{Invent. Math.} \textbf{16} (1975), 181-202.

\bibitem[BK83]{bk83}M. Brin and A. Katok,  On local entropy,  Geometric dynamics (Rio de Janeiro),  \emph {Lecture Notes in Mathematics},  Springer, Berlin,  \textbf{1007} (1983), 30-38.



\bibitem[CL21]{cl21} D. Cheng and Z. Li, Scaled pressure of continuous flows, \emph {Nonlinearity} \textbf{34} (2021), 7829-7841.

\bibitem[CLS21]{cls21} D. Cheng, Z. Li and B. Selmi, Upper metric mean dimensions with potential on subsets, \emph {Nonlinearity} \textbf{34} (2021), 852-867.

\bibitem[CL22]{cl22} D. Cheng and Z. Li, Upper metric mean dimensions for impulsive semi-flows, \emph {J. Diff. Equ.} \textbf{311} (2022), 81-97.

\bibitem[FH12]{fh12} D. J. Feng and W. Huang, Variational principles for topological entropies of subsets, \emph {J. Funct. Anal.} \textbf{263} (2012), 2228-2254.

\bibitem[Gro99]{gromov} M. Gromov,  Topological invariants of dynamical systems and spaces of holomorphic maps: I, \emph{Math. Phys, Anal. Geom.} \textbf{4} (1999), 323-415.

\bibitem[Gut15]{g15} Y. Gutman, Mean dimension and Jaworski-type theorems, \emph{Proc. Lond. Math. Soc.} \textbf{111} (2015),  831-850.


\bibitem[Gut17]{g17} Y. Gutman, Embedding topological dynamical systems with periodic points in cubical shifts, \emph{Ergodic Theory Dynam. Syst.} \textbf{37} (2017), 512-538.

\bibitem[GJ20]{gj20} Y. Gutman and L. Jin,  Mean dimension and an embedding theorem for real flows, \emph{Fund. Math.} \textbf{251} (2020), 161-181.

\bibitem[GLT16]{glt16} Y. Gutman, E. Lindenstrauss and M. Tsukamoto, Mean dimension of $\mathbb{Z}^{k}$ actions, \emph{Geom. Funct. Anal.} \textbf{26} (2016),  778-817.

\bibitem[GS21]{gs20} Y. Gutman and A. \'Spiewak,  Around the variational principle for metric mean dimension,  \emph {Studia Math.} \textbf{261} (2021), 345-360.

\bibitem[GT20]{gt20}  Y. Gutman and M. Tsukamoto, Embedding minimal dynamical systems into Hilbert cubes, \emph {Invent. Math.} \textbf{221} (2020), 113-166.

\bibitem[Kat80]{k80} A. Katok,  Lyapunov exponents, entropy and periodic orbits for diffeomorphisms, \emph{Publ. Math. Inst. Hautes Études Sci.} \textbf{51} (1980), 137-173.

\bibitem[LW00]{lw00} E. Lindenstrauss and B. Weiss,  Mean topological dimension, \emph{Israel J. Math.} \textbf{115} (2000), 1-24. 

\bibitem[LT18]{lt18} E. Lindenstrauss  and M. Tsukamoto, From rate distortion theory to metric mean dimension: variational principle, \emph{IEEE Trans. Inform. Theory} \textbf{64} (2018), 3590-3609.

\bibitem[Pes97]{p97} Y.B. Pesin, Dimension theory in dynamical systems, University of Chicago Press, 1997.

\bibitem[Shi22]{shi} R. Shi, On variational principles for metric mean dimension, to appear in \emph{IEEE Trans. Inform. Theory}, \textbf{68} (2022), 4282-4288.
 

\bibitem[VV17]{vv17}  A. Velozo and R. Velozo, Rate distortion theory, metric mean dimension and measure theoretic entropy, arXiv:1707.05762.

\bibitem[Wal82]{w82} P. Walter, An introduction to ergodic theory,  Springer-Verlag, New York, 1982.

\bibitem[W21]{w21} T. Wang, Variational relations for metric mean dimension and rate distortion dimension, \emph{Discrete Contin. Dyn. Syst.} \textbf{27} (2021),  4593-4608.

\bibitem[Wu21]{wu21}  W. Wu, On relative metric mean dimension with potential and variational principles,  \emph{ J. Dynam.  Diff. Equ.} \textbf{34} (2022),  2313-2335.

\bibitem[YCZ22]{ycz22}  R. Yang, E. Chen and X. Zhou,   Bowen's  equations for upper metric mean dimension with potential, \emph{Nonlinearity} \textbf{35} (2022),  4905-4938.

\bibitem[YZ07]{yz07} X. Ye and G. Zhang,   Entropy points and applications, \emph{Trans. Amer. Math. Soc.} \textbf{359} (2007), 6167-6186.

\end{thebibliography}

\documentclass[a4paper,12pt]{amsart}
\usepackage{amsmath,amsthm}
\usepackage{amsfonts,amsmath,amsthm, amssymb, mathrsfs}
\usepackage{float}
\usepackage{euscript,verbatim}
\usepackage{graphicx}
\usepackage[usenames]{color}
\usepackage[colorlinks,linkcolor=red,anchorcolor=blue,citecolor=blue]{hyperref}
\usepackage{amsmath}
\usepackage{amsthm}
\usepackage[all]{xy}
\usepackage{graphicx} 
\usepackage{amssymb} 
\usepackage{enumerate}
\usepackage{lipsum}
\usepackage{enumitem}

\newtheorem{thm}{Theorem}[section]
\newtheorem{prop}[thm]{Proposition}
\newtheorem{df}[thm]{Definition}
\newtheorem{lem}[thm]{Lemma}
\newtheorem{cor}[thm]{Corollary}

\def\diam{\text{\rm diam}}

\def\over{{ \rm \overline{mdim}}_M(\phi,X,d)}
\def\under{ {\rm\underline{mdim}}_M(\phi,X,d)}
\def \logf{\log \frac{1}{\epsilon}}
\def\ov{\overline{\rm mdim}_M(\phi_1,X,d)}

\def\i{\mu \in \mathcal{M}_{\phi}(X)}
\def\e{\mu \in \mathcal{E}_{\phi}(X)}

\def\ergodic{\mu \in \mathcal{E}_{\phi_1}(X)}
\makeatletter 
\@addtoreset{equation}{section}
\numberwithin{equation}{section}

\title{Metric mean dimension of flows}

\author{Rui Yang, Ercai Chen and Xiaoyao Zhou*
}
\address
{1.School of Mathematical Sciences and Institute of Mathematics, Nanjing Normal University, Nanjing 210023, Jiangsu, P.R.China}
\email{zkyangrui2015@163.com}
\email{ecchen@njnu.edu.cn}
\email{zhouxiaoyaodeyouxian@126.com}
\date{}
\begin{document}
\thispagestyle{empty}
\renewcommand{\thefootnote}{}
\footnote{2020 \emph{Mathematics Subject Classification}:37C45, 37D35}
\footnotetext{\emph{Key words and phrases}:  Continuous flow; Metric mean dimension; Variational principle.}
\footnote{*corresponding author}

\begin{abstract}
The present paper aims to investigate the  metric mean dimension theory of  continuous flows.  We introduce  the notion of metric mean dimension  for continuous flows to characterize the complexity of flows with infinite topological entropy.   For continuous flows,  we establish variational principles for metric mean dimension in terms of  local $\epsilon$-entropy function  and  Brin-Katok $\epsilon$-entropy; For  a class of special  flow, called uniformly Lipschitz  flow, we  establish variational principles for metric mean dimension in terms of Kolmogorov-Sinai $\epsilon$-entropy, Brin-Katok's $\epsilon$-entropy and  Katok's $\epsilon$-entropy. 
\end{abstract}
\maketitle

\section{Introduction}

By a pair $(X,\phi)$ we mean  a   \emph{continuous flow}, where   $X$ is a   compact metrizable  topological space  $X$ with a metric  $d$, $ \phi: X \times \mathbb{R} \rightarrow X$  is   a continuous mapping so that $\phi_{t+s}=\phi_t\circ \phi_s$ for all $t,s \in \mathbb{R}$ and $\phi_t(x):=\phi(x,t)$ denotes  the  homeomorphism on $X$. Let $\mathcal{M}(X)$ denote the set of  Borel probability measures on $X$.   Given  $t \in \mathbb{R}$,  for discrete topological dynamical system $(X,\phi_t)$ the sets of \emph{$\phi_t$-invariant}, \emph{$\phi_t$-ergodic} Borel probability measures on $X$ are denoted by  $\mathcal{M}_{\phi_t}(X)$, $\mathcal{E}_{\phi_t}(X)$, respectively.   A Borel probability  measure $\mu$  on $X$  is said to be \emph{$\phi$-invariant} if $\mu$ is $\phi_t$-invariant for all $t\in \mathbb{R}$. A $\phi$-invariant measure  is said to be \emph{$\phi$-ergodic} if  any  Borel measurable set $B$ with  $\phi_t(B)=B$  for all $t\in \mathbb{R}$  has  measure 0 or 1.  By $\mathcal{M}_{\phi}(X), \mathcal{E}_{\phi}(X)$  we denote the sets of   $\phi$-invariant probability measures, $\phi$-ergodic probability measures, respectively.

In 1999, Gromov \cite{gromov}  introduced   a new topological invariant called  Mean  dimension  for topological dynamical systems. Since then, mean dimension  has confirmed  a powerful tool  solving the embedding problems of dynamical systems \cite{lw00,g15,glt16,g17,gt20}. Later,  Lindenstrauss and Weiss \cite{lw00} introduced  metric mean dimension,  and showed  that metric mean dimension is an upper bound of mean dimension.   It turns out that metric mean dimension is a useful quantity to  characterize the topological complexity of infinite entropy systems.  As the classical variational principle  \cite{w82}  shown,  it is variational principle that  bridges the  ergodic theory and topological dynamics. An important question is how to inject ergodic theoretic ideas into mean dimension theory  by establishing  some new variational principles for metric mean dimension.  In 2018,  Lindenstrauss and Tsukamoto's \cite{lt18} pioneering work shows that there exists variational principle for metric mean dimension  in terms of rate-distortion functions that comes  from information theory.  Changing the candidate rate-distortion functions, Lindenstrauss-Tsukamoto variational principles are still valid \cite{vv17,gs20,shi,wu21}.  
 
The present   paper focus on metric mean dimension  theory of  continuous flows.  On the one hand, there exists intrinsical differences  between the ergodic theory of flow and its discrete samples. For instances, in general an invariant  probability measure   for time one map is not invariant for flow,  and an ergodic probability  measure  for flow is  not necessarily ergodic for time one map. Hence one can not directly derive  variational principle for metric mean dimension of flows based the previous work.  On the other hand,  Abramov entropy formulas \cite{a59} show that   both topological entropy and measure-theoretic entropy  of  discrete samples $(\phi_t)_{t\in \mathbb{R}}$ are   equal to the absolute value of $t$ times their corresponding  entropies of $\phi_1$, yet for  different discrete samples of flow,   the  metric mean dimensions of the phase space and the measure-theoretic metric mean dimensions of invariant measures (defined by measure-theoretic $\epsilon$-entropies) may allow different ``speeds" to  approximate the infinite entropy along with the time. The  two obstacles lead to   some  significant difficulties when establishing variational principles for  metric mean dimension of  continuous flows.  
  
To overcome the first obstacle, we  introduce the notion of local $\epsilon$-entropy  function  of flows inspired by  \cite{yz07,shi} and establish a variational principle for metric mean dimension in terms of the local $\epsilon$-entropy  function of flows.  
 \begin{thm}\label{thm 1.1}
 Let $(X,\phi)$ be a  continuous  flow with a metric $d$. Then 
 \begin{align*}
 \overline{\rm mdim}_M(\phi,X, d)&=\limsup_{\epsilon \to 0}\frac{1}{\log \frac{1}{\epsilon}} \sup_{x\in X} h_d (x,\epsilon,\phi)\\
 \underline{\rm mdim}_M(\phi,X, d)&=\liminf_{\epsilon \to 0}\frac{1}{\log \frac{1}{\epsilon}} \sup_{x\in X} h_d (x,\epsilon,\phi).
 \end{align*}
  \end{thm}
  
Besides,  motivated by the work \cite{b73,fh12,w21,cls21,ycz22} we introduce the notion of  Bowen  metric mean dimension  on  subsets of continuous flows, which  allows us to establish variational principle for Bowen metric mean dimension for compact subsets  in terms of  Brin-Katok  $\epsilon$-entropy and then extend the variational principles  to  the whole phase space. 
    
\begin{thm}\label{thm 1.2}
Let $(X,\phi)$ be a  continuous  flow with a metric $d$. Then 
\begin{align*}
\overline{\rm mdim}_M(\phi,X, d)
&=\limsup_{\epsilon \to 0}\frac{1}{\log \frac{1}{\epsilon}} \sup_{\mu \in \mathcal{M}(X)}\underline{h}_\mu^{BK}(\phi,\epsilon)\\
&=\limsup_{\epsilon \to 0}\frac{1}{\log \frac{1}{\epsilon}} \sup_{\mu \in \mathcal{M}(X)}\overline{h}_\mu^{BK}(\phi,\epsilon),
\end{align*}
The result is also true for $\underline{\rm mdim}_M(\phi,X, d)$ by changing $\limsup_{\epsilon \to 0}$ into $\liminf_{\epsilon \to 0}$.
 \end{thm}
 
 To overcome the second obstacle,  we   need an  auxiliary condition on the flow to offset the differences  in the  topological and measure-theoretic aspects  for  continuous flows and its discrete samples. In  this case,  we can establish variational principles for metric mean dimension of a  class  of special  continuous flow called  uniformly Lipschitz flows.

\begin{thm}\label{thm 1.3}
Let $(X,\phi)$ be a uniformly Lipschitz flow with a metric $d$.  Then for every $F(\mu,\epsilon)\in \mathcal{D}$
\begin{align*}
\over&=\limsup_{\epsilon \to 0}\frac{1}{\logf}\sup_{\mu \in \mathcal{M}_{\phi}(X)}F(\mu,\epsilon)\\
&=\limsup_{\epsilon \to 0}\frac{1}{\logf}\sup_{\mu \in \mathcal{E}_{\phi}(X)}F(\mu,\epsilon),
\end{align*}
where $F(\mu,\epsilon)$ is chosen from  the candidate set
$$\mathcal{D}=\left\{\inf_{\diam (P) \leq \epsilon}\limits h_\mu(\phi_1, {P}), 
\overline{h}_\mu^{BK}(\phi_1,\epsilon),\underline{h}_{\mu}^K(\epsilon, \delta,\phi_1),\overline{h}_{\mu}^K(\phi_1,\epsilon, \delta)\right\}.$$
The result is also true for $\underline{\rm mdim}_M(\phi,X, d)$ by changing $\limsup_{\epsilon \to 0}$ into $\liminf_{\epsilon \to 0}$.
 \end{thm}

 We  remark that   if the scale function $S=\logf$ and the potential function $f=0$, then   the condition A  can be  removed in  \cite[Theorem C]{cl21} and the supremum can only  range over all  ergodic measures of $\phi$ by Theorem \ref{thm 1.3}.

The rest of this paper is organized as follows. In section 2,  we introduce the notions of  the metric mean dimension of flows  and derive some elementary properties.  In section 3,  we  give the proof of  Theorems \ref{thm 1.1}, \ref{thm 1.2} and  \ref{thm 1.3}.

\section{Preliminary}


In this section, we  introduce the notion of metric mean dimension  of continuous flows defined by spanning sets and separated sets,  and derive some elementary properties related with metric mean dimension, including the classical Lindenstrauss-Weiss inequality and Abramov type  formula of metric mean dimension.

 Let  $t\in \mathbb{R}$, $n\in \mathbb{N}$ and $x,y \in X$.  The \emph{$t$-th Bowen metric  for flow $\phi$}, \emph{$n$-th Bowen metric for time $t$-map $\phi_t$}  are respectively given by 
\begin{align*}
d_t(x,y):&=\max_{s\in [0,t]} d(\phi_s x, \phi_s y),\\
d_{n,\phi_t}(x,y):&=\max_{ j\in \{0,...,n-1\}} d(\phi_{tj}x,\phi_{tj} y).
\end{align*}
Then the  $(t,\epsilon,\phi)$-ball of $x$ and   the  $(n,\epsilon,\phi_t)$-ball of $x$  are respectively defined  by  
\begin{align*}
B_t(x,\epsilon,\phi)&=\{y\in X: d_t(x,y)<\epsilon\},\\
B_n(x,\epsilon,\phi_t)&=\{y\in X: d_{n,\phi_t}(x,y)<\epsilon\}.
\end{align*}
Clearly,  both sets $B_t(x,\epsilon,\phi)$ and  $B_n(x,\epsilon,\phi_t)$ are open due to the continuity of $\phi$.

Fix  a non-empty subset $Z\subset X$ and $\epsilon >0$. A set $E\subset X$ is  \emph{a $(t,\epsilon)$-spanning set of $Z$} if  for any $x \in Z$, there  exists  $y\in E$ such that $d_t(x,y)<\epsilon.$ The smallest cardinality of $(t,\epsilon)$-spanning set of $Z$  is denoted by $r_t(\phi,Z,d,\epsilon)$. A set $F\subset Z$ is  \emph{a $(t,\epsilon)$-separated set of $Z$} if $d_t(x,y)\geq\epsilon$ for any $x,y \in F$ with $x\not= y$. The  largest  cardinality  of $(t,\epsilon)$-separated set of $Z$  is denoted by $s_t(\phi,Z, d,\epsilon)$.  Put
$$r(\phi, Z,d,\epsilon)=\limsup_{t\to \infty} \frac{1}{t} \log r_t(\phi,Z,d,\epsilon)$$
and
$$s(\phi, Z,d,\epsilon)=\limsup_{t\to \infty} \frac{1}{t} \log s_t(\phi,Z,d,\epsilon).$$
 By a standard method \cite{w82}, we  have $r(\phi, X,d,\epsilon)\leq s(\phi, X,d,\epsilon) \leq r(\phi, X,d,\frac{\epsilon}{2}).$

\begin{df}
Let $(X,\varphi)$ be a  continuous flow  with a metric $d$. The  upper  and lower metric mean dimensions of  $X$  for flow  $\phi$  are defined by 
\begin{align*}
\over&=\limsup_{\epsilon\to 0}\frac{r(\phi,X,d,\epsilon)}{\log \frac{1}{\epsilon}}=\limsup_{\epsilon\to 0}\frac{s(\phi,X,d,\epsilon)}{\log \frac{1}{\epsilon}},\\
\under&=\liminf_{\epsilon\to 0}\frac{r(\phi,X,d,\epsilon)}{\log \frac{1}{\epsilon}}=\liminf_{\epsilon\to 0}\frac{s(\phi,X,d,\epsilon)}{\log \frac{1}{\epsilon}}.\\
\end{align*}
\end{df}
For time $t$-map $\phi_t$, using $d_{n,\phi_t}$ metric one can similarly define  the quantities $r(\phi_t, X,d,\epsilon),$ $ s(\phi_t, X,d,\epsilon)$, and    upper and lower metric mean dimensions  $\overline{\rm{mdim}}_M(\phi_t, X,d)$,  $\underline{\rm{mdim}}_M(\phi_t, X,d)$, respectively.

Obviously, metric mean dimension for flows  depends on the metric on $X$ and hence are not topological invariant.  Recall that Bowen and Ruelle \cite{br75}  defined  topological entropy   of   continuous flows as 
$$h_{top}(\phi,X)=\lim\limits_{\epsilon \to 0}r(\phi, X,d,\epsilon)=\sup_{\epsilon>0}r(\phi, X,d,\epsilon).$$
Notice that  the continuous flows have zero  metric mean dimension  if its topological entropy is finite. Hence  metric mean dimension  for flows is a useful quantity to characterize  flow admitting   infinite topological entropy. 

We  proceed to  derive some basic properties of  metric mean dimension of continuous flows.
 \begin{prop}
Let $(X,\phi)$ be a continuous flow with a metric $d$. Then  for every $\tau >0$
 \begin{align*}
\over=\limsup_{\epsilon\to 0}\limsup_{n\to \infty}\frac{r_{n\tau}(\phi,X,d,\epsilon)}{n\tau\log\frac{1}{\epsilon}}\\
\under=\liminf_{\epsilon\to 0}\limsup_{n\to \infty}\frac{r_{n\tau}(\phi,X,d,\epsilon)}{n\tau \log\frac{1}{\epsilon}}.
 \end{align*}
\end{prop}

A flow naturally induces a  discrete topological dynamical system $(X,\phi_t),t\in \mathbb{R}.$ The following proposition  is an analogue of  Abramov measure-theoretic entropy formula \cite{a59}, which examines the relationship   between  the metric mean dimension of a flow and its discrete samples. The following  argument is also used  by Li and Cheng \cite[Defintiion 4.1]{cl21} to study the scaled pressure of  fixed-point free  continuous flows. 
\begin{df}\label{def 2.3}
A  continuous flow $(X,\phi)$ is said to be a  uniformly Lipschitz flow if for any $t_0>0$,  there exists $L(t_0)>0$ such that for any  $\epsilon>0$ and $x,y\in X$ with $d(x,y)\leq \frac{\epsilon}{L(t_0)}$, one has $$d(\phi_t(x),\phi_t(y))<\epsilon$$
for all $s\in [0,t_0]$. 
\end{df}

\begin{prop}\label{prop2.4}
Let $(X,\phi)$ be a continuous flow with a metric $d$.  Then for every $\tau\in \mathbb{R_+} \backslash \{0\}$
\begin{align*}
\frac{1}{\tau}\overline{\rm mdim}_M(\phi_{\tau},X, d) &\leq \over,\\
\frac{1}{\tau}\underline{\rm mdim}_M(\phi_{\tau},X, d) &\leq \under,
\end{align*}
and the equalities hold  if $(X,\phi)$ is a  uniformly Lipschitz flow.
\end{prop} 

\begin{proof}
Fix $\tau >0$ and $\epsilon >0$. Choose a subsequence $n_k$ as $k\to \infty$  that converges to $\infty$  and satisfies that
$\lim_{k\to \infty}\frac {\log r_{n_k}(\phi_{\tau}, X, d, \epsilon)}{n_k}=r(\phi_{\tau}, X, d, \epsilon)$.
For every $k$,  choose  a subsequence $t_k$ such that 
$n_k\tau \leq t_k <(n_k+1)\tau.$
Since  $B_{t_k}(x,\epsilon,\phi)\subset B_{n_k}(x,\epsilon,\phi_{\tau})$, we have  $$\frac{r_{n_k}(\phi_{\tau},X,d,\epsilon)}{n_k}\leq  \frac{r_{t_k}(\phi,X,d,\epsilon)}{t_k}\cdot\frac{t_k}{n_k}.$$
Noticing that  $\lim_{k\to \infty}\frac{t_k}{n_k}=\tau$,  we get
$ \frac{r(\phi_{\tau}, X, d, \epsilon)}{\tau}\leq  r(\phi,X,d,\epsilon)$ by letting $k  \to \infty$, which implies the desired results. 

Assuming  that $(X,\phi)$ is a  uniformly Lipschitz flow, then  there is  $L(\tau)>0$ so that $d(\phi_{s}x, \phi_{s}y)<\epsilon$ for all $0\leq s \leq \tau$  if  $x,y \in X$ with $d(x,y)<\frac{\epsilon}{L(\tau)}$.  Hence $B_n(x,\frac{\epsilon}{L(\tau)},\phi_{\tau})\subset B_{n\tau}(x,\epsilon,\phi)$  for every $n\in \mathbb{N}$.  Using the fact that
$r(\phi,X,d,\epsilon)=\limsup_{n\to \infty}\limits\frac{r_{n\tau}(\phi,X,d,\epsilon)}{n\tau},$ 
we obtain that 
$$\frac{r(\phi_{\tau},X,d,\frac{\epsilon}{L(\tau)})}{\tau}\geq  r(\phi,X,d,\epsilon).$$
This completes the proof.
\end{proof}

We  interpret  why  uniformly Lipschitz flow is needed for obtaining the converse inequalities.   Given continuous flow $(X,\phi)$, by the continuity of $\phi$, there exists $0<\delta(\epsilon)<\epsilon$ such that $d(x,y)<\delta(\epsilon)$ so that  $d(\phi_s x,\phi_s y)<\epsilon$ for all $s\in [0,\tau]$. Similarly, we have 
$\frac{r(\phi_{\tau}, X, d,\delta(\epsilon))}{\tau}\geq  r(\phi,X,d,\epsilon).$
According to the definition of metric mean dimension, one  can formulate  the following inequality
$$\frac{\log \frac{1}{\delta(\epsilon)}}{\log \frac{1}{\epsilon}}\cdot \frac{r(\phi_{\tau}, X,d,\delta(\epsilon))}{\tau \log \frac{1}{\delta(\epsilon)}}\geq \frac{r(\phi,X,d,\epsilon)}{\logf}.$$
However, we fail to determine  whether $\limsup_{\epsilon\to 0}\frac{\log \frac{1}{\delta(\epsilon)}}{\log \frac{1}{\epsilon}} =1$ or not.

Recall that the \emph{mean dimension} of a  continuous flow $(X,\phi)$ introduced by Gutman and  Jin \cite{gj20} is given by 
$$\text{mdim}(X,\phi)=\lim_{\epsilon\to 0}\lim_{n\to \infty}\frac{\text{Wdim}_{\epsilon}(X,d_n)}{n},$$
where ${\rm Wdim}_{\epsilon}(X,d_n)$ is  defined by $(d_n,\epsilon)$-embedding mapping. Replacing $d_n$  by  the metric $d_{n,\phi_1}$,  it reduces to the classical notion  $\text{mdim}(X,\phi_1)$ for time one-map  $\phi_1$.  Invoking the classical  Lindenstrauss-Weiss's  inequality  and Proposition 2.4,  we   can  extend  the classical  inequality to continuous flows.  
\begin{cor}
Let $(X,\phi)$ be a continuous flow with a metric $d$. Then 
\begin{align*}
\text{\rm mdim}(X,\phi)\leq \under \leq \over.
\end{align*}

\end{cor}
\begin{proof}
By \cite[Proposition 2.5]{gj20},  we have  $$\text{mdim}(X,\phi)=\text{mdim}(X,\phi_1).$$
Together with the   fact   $\text{mdim}(X,\phi_1)\leq \text{\underline{\rm mdim}}_M(X,\phi_1,d)$ \cite[Theorem 4.2]{lw00} and Proposition 2.4,  we  get the desired result. 

\end{proof} 
\section{Proof of main results}
\subsection{Proof of Theorem  \ref{thm 1.1}}
In this subsection, we   give the proof of  Theorem  \ref{thm 1.1}.

In \cite{yz07}, Ye and Zhang  introduced the  notion of  local entropy function  to study the uniform entropy points and proved  that $$h_{top}(\phi_1,X)=\sup_{x\in X} {h(\phi_1,x)},$$ where  $h_{top}(\phi_1,X)$ and $h(\phi_1,x)$ denote  topological entropy of $X$, the local entropy function at $x$  for time  one map, respectively. We extend this notion to continuous flows to establish an analogous  variational principle for  metric mean dimension. 

Given  $\epsilon>0$ and $x\in X$,  we define the \emph{local $\epsilon$-entropy function  at $x$ with respect to $\phi$} as
$$h_d(x, \epsilon,\phi)=\inf \{r(\phi, K,d,\epsilon): K~\text{is a closed neighborhood of}~x \}.$$

\begin{prop} \label{prop 2.1}
Let $(X,\phi)$ be a continuous flow with a metric $d$.  Suppose that $X$ is a finite union of closed subset  $K_i$, $i=1,...,N$. Then for every $\epsilon>0$, 
$$r(\phi,X,d,\epsilon)=\max_{1 \le j \le N} r(\phi,K_j,d,\epsilon).$$
Consequently,
$\over= \max_{1 \leq j \le N} \limits \overline{\rm mdim}_M(\phi,K_j,d).$   
\end{prop}

\begin{proof}
Fix $\epsilon>0$.  Clearly, one has $r(\phi,X,d,\epsilon)\geq\max_{1 \le j \le N} r(\phi,K_j,d,\epsilon).$  For each $t>0$, there exists $ j_{(\epsilon,t)}\in \{1,...,N\}$ such  that $r_t(\phi,K_{j_{(\epsilon,t)}},d,\epsilon)=$\\
$\max_{1 \le j \le N}$
$ r_t(\phi,K_j,d,\epsilon)$. Choose  a subsequence $t_k$ that converges to $\infty$  as $k \to \infty$ so that  
$$r(\phi,X,d,\epsilon)=\lim_{k\to \infty}\frac{r_{t_k}(\phi,X,d,\epsilon)}{t_k}$$
and
$$r_{t_k}(\phi,K_{j_{(\epsilon)}},d,\epsilon)=\max_{1 \le j \le N} r_{t_k}(\phi,K_j,d,\epsilon)$$
for all $k\in \mathbb{N}$, where  $j_{\epsilon } \in \{1,...,N\}$ only depends on $\epsilon$.
This yields that 
$$r(\phi,X,d,\epsilon)\leq r(\phi,K_{j_{(\epsilon)}},d,\epsilon)\leq \max_{1 \le j \le N} r(\phi,K_j,d,\epsilon).$$ 
\end{proof}

By means of  Proposition \ref{prop 2.1}, we give the proof of Theorem  \ref{thm 1.1}.
\begin{proof}[Proof of Theorem \ref{thm 1.1}]
Fix  $\epsilon >0$. The inequality
$\sup_{x\in X}h_d(x,\epsilon,\phi)\leq r(\phi,X,\epsilon)$ is clear. 

Cover $X$ with the closed ball family $\{\bar{B}_1^1,\bar{B}_2^1,...,\bar{B}_{n_1}^1\}$ whose diameter is  $1$. By  Proposition \ref{prop 2.1},  there exists $1\leq j_1\leq n_1$ such that
$r(\phi,X,d,\epsilon)=r(\phi,\bar{B}_{j_1}^1,d,\epsilon).$
Cover the closed ball $\bar{B}_{j_1}^1$ with closed ball  family $\{\bar{B}_1^2,
\bar{B}_2^2,...,\bar{B}_{n_2}^2\}$, where  the diameter of   every subset  $\bar{B}_i^2$ of $\bar{B}_{j_1}^1$  is  $\frac{1}{2}$. Applying  Proposition \ref{prop 2.1} again,  there exists  $1\leq j_2\leq n_2$ such that 
$r(\phi,X,d,\epsilon)=r(\phi,\bar{B}_{j_2}^2,d,\epsilon).$ Proceeding this procedure,
for every $k\geq 1$  there is a closed ball $\bar{B}_{j_k}^k$ with diameter at most $\frac{1}{k}$ such that
$r(\phi,X,d,\epsilon)=r(\phi,\bar{B}_{j_k}^k,d,\epsilon).$ Set $\mathop\cap_{k\geq 1}\limits\bar{B}_{j_k}^k=\{x_0\}$. Then for any closed neighborhood $K$ of $x_0$ we can choose  sufficiently large $k_0\in \mathbb{N}$ such that 
$\bar{B}_{j_{k_0}}^{{k_0}}\subset K$. So $r(\phi,X,d,\epsilon)\leq r(\phi,K,d,\epsilon)$, which implies that
$$r(\phi,X,d,\epsilon)\leq h_d(x_0,\epsilon,\phi)\leq \sup_{x\in X}h_d(x,\epsilon,\phi).$$
This completes the proof.
\end{proof}

\subsection{Proof of Theorem \ref{thm 1.2}} 
 In this subsection, we  introduce the notion of upper and lower Brin-Katok  $\epsilon$-entropies  of $\phi$, and  give the proof  Theorem  \ref{thm 1.2}.
  
Given $\mu \in \mathcal{M}(X)$ and $\epsilon >0$, we   define \emph{upper and lower Brin-Katok  $\epsilon$-entropies  of $\phi$} as
\begin{align*}
\overline{h}_{\mu}^{BK}(\phi, \epsilon)&=\int \limsup_{t\to \infty}-\frac{\log \mu (B_t(x,\epsilon,\phi))}{t}d\mu,\\
\underline{h}_{\mu}^{BK}(\phi, \epsilon)&=\int \liminf_{t\to \infty}-\frac{\log \mu (B_t(x,\epsilon,\phi))}{t}d\mu.
\end{align*} 

We  first define  Bowen metric mean dimension by means of  Carath\'eodory-Pesin structure \cite{p97}, which allows us to  borrow some tools from geometric measure theory to  obtain the variational relation between metric mean dimension and   Brin-Katok  $\epsilon$-entropy.

\begin{df}\label{df 3.2}
Let  $Z\subset X$ be a non-empty subset, $\epsilon >0, N\in\mathbb{N}$, and $s\geq 0$. Put 
$$M_{N.\epsilon}^{s}(\phi,Z,d)=\inf\{\sum_{i\in I}\limits  e^{-n_i s}\},$$
where the infimum  is taken over all  finite or countable   covers $\{B_{n_i}(x_i,\epsilon,\phi)\}_{i\in I}$ of $Z$ with $n_i \geq N.$

Since $M_{N.\epsilon}^{s}(\phi,Z,d)$ is non-decreasing when $N$ increases, so the limit
$M_{\epsilon}^s(\phi,Z,d)=\lim\limits_{N\to \infty}M_{N.\epsilon}^{s}(\phi,Z,d)$ exists.  There is a critical value  of parameter $s$ for $M_{\epsilon}^s(\phi,Z,d)$  jumping from $\infty$ to $0$, which  is denoted by 
\begin{align*}
h_{top}^B(\phi,Z,d,\epsilon)&=\inf\{s:M_{\epsilon}^s(\phi,Z,d)=0\}\\
&=\sup\{s:M_{\epsilon}^s(\phi,Z,d)=\infty\}.
\end{align*}
We  define Bowen upper metric mean dimension of $\phi$  on the set $Z$ as 
\begin{align*}
\overline{\rm mdim}_M^B(\phi,Z,d)&=\limsup_{\epsilon \to 0}\frac{h_{top}^B(\phi,Z,d,\epsilon)}{\log \frac{1}{\epsilon}}.
\end{align*}
\end{df}


When $Z=X$, we adapt Bowen's approach \cite{b73} to show  metric mean dimensions of  $X$  defined by   Carath\'eodory-Pesin structure and spanning set are equivalent.

\begin{prop} \label{prop 2.9}

 Let  $(X,\phi)$ be a  continuous flow with a metric $d$. Then 
 \begin{align*}
\overline{\rm {mdim}}_{M}(\phi,X,d)&=\overline{\rm mdim}_M^B(\phi,X,d),\\
\underline{\rm {mdim}}_{M}(\phi,X,d)&=\underline{\rm mdim}_M^B(\phi,X,d).
 \end{align*}

\end{prop}
 
\begin{proof}
It  suffices to show the first equality and the second one can be proved  in similar manner.
The inequality $\overline{\rm mdim}_M^B(\phi,X)\leq \overline{\rm {mdim}}_{M}(\phi,X)  $  follows by using the fact every   $(n,\epsilon)$- spanning set  of $X$ with the  minimal cardinality $r_n(\phi,X,d,\epsilon)$  also  covers  $X$.

Fix $\epsilon>0$ and let  $s>h_{top}^B(\phi,Z,d,\epsilon)$. By the  compactness of $X$, there are  $N_0$  and a finite open cover $\{B_{t_i}(x_i,\epsilon,\phi)\}_{i\in I}$ of $X$  with $t_i \geq N_0$ and 
$\sum_{i\in I} {e}^{-st_i}<1$, where $I$ is finite index set. It follows that
\begin{align*}
\sum_{k=1}^{\infty}\sum_{j_1,...,j_k\in I}e^{-s(t_{j_1}+\cdots t_{j_k})}= \sum_{k=1}^{\infty}(\sum_{i\in  I}e^{-st_i})^k <\infty.
\end{align*}

Let $M=\max_{i\in I} t_i$. We  define $c_0=0$ and $c_i=\sum_{s=0}^{k-1}t_{j_k},j_k\in I $ for every $k \geq 1$.  For sufficiently large  $N \geq N_0$,  consider the family
\begin{align*}
\mathcal{F}_N:=\{\mathop\cap_{i=0}^{k}\limits \phi_{-c_i}B_{t_{j_i}}(x_{t_{j_i}},\epsilon,\phi): k\geq 0, 
N\leq  \sum_{i=0}^{k}t_{j_i}<N+M\}.
\end{align*} 
Without loss of generality, assuming that each  $A\in \mathcal{F}_N$ is not empty,  then we choose $x_A\in A $ so that  $A\subset B_N(x_{A},2\epsilon,\phi).$ This implies that 
$X\subset \mathop\cup_{A\in \mathcal{F}_N}\limits B_N(x_{A},2\epsilon,\phi).$  Therefore,
\begin{align*}
r_N(\phi,X,d,2\epsilon)e^{-sN}&\leq \# \mathcal{F}_N\cdot e^{-sN}\\
&\leq e^{sM} \sum_{k\geq0}\sum_{N\leq t_{j_1}+...+t_{j_k}<N+M} e^{-s(t_{j_1}+\cdots t_{j_k})}   <\infty.
\end{align*}

This implies that  $r(\phi,X,d,2\epsilon)\leq s$.  We get $r(\phi,X,d,2\epsilon)\leq h_{top}^B(\phi,Z,d,\epsilon)$ after  letting $s \to h_{top}^B(\phi,Z,d,\epsilon)$.
\end{proof}

\begin{df}
Let $f: X \rightarrow \mathbb{R}$ be a bounded function, and let  $N\in \mathbb{N}$, $\epsilon>0,  s \geq 0$.  Put
$$W_{N,\epsilon}^s(\phi,f,d)=\inf\{\sum_{i\in I}\limits  c_ie^{-n_i s}\},$$
where the infimum  is taken over all  finite or countable families  $\{B_{n_i}(x_i,\epsilon,\phi)\}_{i\in I}$ with  $0<c_i<\infty$, $x_i\in X$  and  $n_i\geq N$,  so that 
$\sum_{i\in I}\limits c_i\chi_{B_{n_i}(x_i,\epsilon,\phi)}\geq f.$ 

Let $Z\subset X$ be a non-empty subset. Set $W_{N,\epsilon}^s(\phi,Z,d):=W_{N,\epsilon}^s(\phi,\chi_Z,d)$  and $W_{\epsilon}^s(\phi,Z,d)=\lim\limits_{N\to \infty}W_{N,\epsilon}^s(\phi,Z,d).$
There is a critical value  of parameter $s$  for $W_{\epsilon}^s(\phi,Z,d)$  jumping from $\infty$ to $0$. 
\begin{align*}
h_{top}^{WB}(\phi,Z,d,\epsilon):&=\inf\{s:W_{\epsilon}^s(\phi,Z,d)=0\}\\
&=\sup\{s:W_{\epsilon}^s(\phi,Z,d)=\infty\}.
\end{align*}

We define  weighted Bowen  upper and  lower metric mean dimension of $\phi$  on the set $Z$ as 
\begin{align*}
\overline{\rm mdim}_M^{WB}(\phi,Z,d)&=\limsup_{\epsilon \to 0}\frac{h_{top}^{WB}(\phi,Z,d,\epsilon)}{\log \frac{1}{\epsilon}},\\
\underline{\rm mdim}_M^{WB}(\phi,Z,d)&=\liminf_{\epsilon \to 0}\frac{h_{top}^{WB}(\phi,Z,d,\epsilon)}{\log \frac{1}{\epsilon}}.
\end{align*}

\end{df}

Analogous to \cite[Proposition 3.2]{fh12},   one can   show that    weighted Bowen metric mean dimension  coincides with Bowen metric mean dimension.

\begin{prop}\label{prop 2.11}
Let $Z\subset X$ be a non-empty subset.  Then for any $s\geq 0, \epsilon >0, \delta >0$ and $N\in \mathbb{N}$, 
$$M_{N,6\epsilon}^{s+\delta}(\phi,Z,d)\leq W_{N,\epsilon}^s(\phi,Z,d)\leq M_{N,\epsilon}^s(\phi,Z,d).$$
Consequently,
\begin{align*}
\overline{\rm{mdim}}_M^B(\phi,Z,d)&=\overline{\rm mdim}_M^{WB}(\phi,Z,d),\\
\underline{\rm{mdim}}_M^B(\phi,Z,d)&=\underline{\rm mdim}_M^{WB}(\phi,Z,d).
\end{align*} 
\end{prop}

Using weighted Bowen metric mean  dimension, inspired by \cite[Lemma 3.4]{fh12} we can  define a positive  and bounded  linear functional to  produce a Borel probability measures by Riesz  representation theorem. The  following  is the  Frostman's lemma for continuous flows.

\begin{lem}\label{lem 2.12}
Let $K$ be a non-empty compact subset of $X$ and $ s \geq0, \epsilon >0, N\in \mathbb{N}$. Set $c:=W_{N,\epsilon}^s(\phi,K,d)>0$. Then there exists a Borel probability measure $\mu \in M(X)$ such that $\mu (K)=1$ and 
for any $x\in X, n\geq N$, 
$$\mu(B_n(x,\epsilon,\phi))\leq \frac{1}{c}e^{-s n}.$$

\end{lem}

The following  theorem gives  a variational principle for  Bowen metric mean dimension  on compact subsets in the context of infinite entropy, which is an analogue of variational principle for Bowen topological entropy\cite{fh12}.
\begin{thm}\label{thm 2.13}
Let $(X,\phi)$ be a continuous flow  with a metric $d$ and $K$ be a non-empty compact subset of $X$. Then  
\begin{align*}
\overline{\rm mdim}_M^B(\phi,K, d)
&=\limsup_{\epsilon \to 0}\frac{1}{\log \frac{1}{\epsilon}} \sup\{\underline{h}_\mu^{BK}(\phi,\epsilon):\mu \in \mathcal{M}(X),\mu(K)=1\},\\
\underline{\rm mdim}_M^B(\phi,K, d)&=\liminf_{\epsilon \to 0}\frac{1}{\log \frac{1}{\epsilon}} \sup\{\underline{h}_\mu^{BK}(\phi,\epsilon):\mu \in \mathcal{M}(X),\mu(K)=1\}.
\end{align*}
\end{thm}
\begin{proof}
Fix $\epsilon >0$ and  $\mu \in \mathcal{M}(X)$ with  $\mu (K)=1$. Assume that $\underline{h}_\mu^{BK}(\phi,\epsilon)>0$ and let $s<\underline{h}_\mu^{BK}(\phi,\epsilon)$.   By a standard method, one  can  find  a Borel set $E\subset K$  and  $N \in \mathbb{N}$ such that $\mu(E)>0$ and $$-\frac{\log\mu(B_n(x,\epsilon,\phi))}{n}>s$$ for any $x\in E$ and $n\geq N$.  Let $\{B_{n_i}(x_i,\frac{\epsilon}{2},\phi)\}_{i\in I}$ be a   finite  or countable  cover of $E$  with $n_i \geq N$. We  assume that  $B_{n_i}(x_i,\frac{\epsilon}{2},\phi)\cap E\not= \emptyset$ for every $i\in I$. 
Choose $y_i \in  B_{n_i}(x_i,\frac{\epsilon}{2},\phi)\cap E$ for each $i \in I$ so that  $\cup_{i\in I}B_{n_i}(y_i,\epsilon,\phi)\supseteq E$. This yields that 
\begin{align*}
\sum_{i\in  I}e^{-sn_i}\geq \sum_{i\in  I}\mu(B_{n_i}(y_i,\epsilon,\phi))\geq \mu(E)>0.
\end{align*}
Therefore, $ h_{top}^B(\phi,K,d,\frac{\epsilon}{2})\geq h_{top}^B(\phi,E,d,\frac{\epsilon}{2})\geq s$. Letting $s \to \underline{h}_\mu^{BK}(\phi,\epsilon)$, we obtain that $\underline{h}_\mu^{BK}(\phi,\epsilon)\leq h_{top}^B(\phi,K,d,\frac{\epsilon}{2})$ for every $\mu \in  \mathcal{M}(X)$ with $\mu(K)=1$.

On  the other hand, let 
$h_{top}^{WB}(\phi,K,d,\epsilon)>0$ and $s<h_{top}^{WB}(\phi,K,d,\epsilon)$. Then there exists $N_0$ such that  $c:= W_{N_0,\epsilon}^s(\phi,K,d)>0$. By Lemma \ref{lem  2.12},   there is  a $\mu \in \mathcal{M}(X)$ so that $\mu(K)=1$ and 
$$\mu(B_n(x,\epsilon,\phi))\leq \frac{1}{c}e^{- ns}$$
for any $x\in K$ and $n\geq N_0$.
It follows that $\underline{h}_\mu^{BK}(\phi,\epsilon)\geq s$. Letting $s\to h_{top}^{WB}(\phi,K,d,\epsilon)$ and  then using  Proposition \ref{prop 2.11},  we know that  $$h_{top}^{B}(\phi,K,d,6\epsilon)\leq \sup\{\underline{h}_\mu^{BK}(\phi,\epsilon):\mu \in \mathcal{M}(X), \mu(K)=1\}.$$
This completes the proof.
\end{proof}

Combining  Theorem \ref{thm 2.13}, we   give the proof of  Theorem  \ref{thm 1.2}.
\begin{proof}[Proof of Theorem \ref{thm 1.2}]
Fix $\epsilon >0$ and $\mu \in \mathcal{M}(X)$ with  $\mu (K)=1$.   Let $s<\overline{h}_\mu^{BK}(\phi,\epsilon)$. By a standard method, there exists  a Borel  subset $E$  of $X$ such that $\mu(E)>0$ and  $ \limsup_{n\to \infty}\limits-\frac{\log \mu (B_n(x,\epsilon,\phi))}{n}>s$ for any  $x\in E$. Put  $E_n=\{x\in E: \mu(B_n(x,\epsilon,\phi)) < e^{-ns}\}$. Then $E=\cup_{n\geq N}E_n$ for every $N\geq 1$. Hence, one can choose $n\geq N$ (depending on $N$) so that $\mu(E_n)\geq \frac{1}{n(n+1)}\mu(E)$ for every $N\geq 1$.  Let  $F_n \subset E_n$ be an 
 $(n,\epsilon)$-separated set  of $E_n$ with the  largest cardinality $s_n(\phi,E_n,d,\epsilon)$. Then
 $$0<\mu(E_n)\leq \sum_{x\in F_n}\mu(B_n(x,\epsilon,\phi))\leq s_n(\phi,E_n,d,\epsilon)\cdot e^{-ns},$$
which shows $s_n(\phi,X,d,\epsilon)\geq  \mu(E_n) e^{ns}$.   So $s\leq s(\phi,X,d,\epsilon)$. Letting $s \to \overline{h}_\mu^{BK}(\phi,\epsilon)$,  we have   $\overline{h}_\mu^{BK}(\phi,\epsilon)\leq  s(\phi,X,d,\epsilon).$ Together with  Proposition \ref{prop 2.9} and Theorem  \ref{thm 2.13},   one has Theorem \ref{thm 1.2}.

\end{proof}

\subsection{Proof of Theorem 1.3}
In this subsection,  we recall that the definitions of  \emph{Kolmogorov-Sinai $\epsilon$-entropy} \cite{gs20}  and \emph{Katok's entropy} \cite{k80} and  give the proof  of  Theorem  \ref{thm 1.3}. Throughout this section, $L:=L(1)$ is given as Definition \ref{def 2.3}. 
\begin{itemize}[leftmargin = 6pt]
\item  \textbf{Kolmogorov-Sinai $\epsilon$-entropy} 
Given a finite (Borel) measurable  partition  $P$ of $X$, we denote by ${\diam}P=\sup_{A\in P}{{\diam}(A,d)}$ the \emph{diameter} of $P$.   Let $t\in \mathbb{R}$,  $\mu \in \mathcal{M}_{\phi}(X)$. Then $\mu$ is also $\phi_t$-invariant.   For time $t$-map,  let $h_\mu(\phi_t)$ denote the measure-theoretic entropy of $P$ w.r.t. $\mu$ \cite{w82}.  

Given $\epsilon>0$,   we  define 
\emph{Kolmogorov-Sinai $\epsilon$-entropy of $\mu$ w.r.t. $\phi_1$} as
$$\inf_{\diam (P) \leq \epsilon}h_\mu(\phi_1, P),$$
where  the infimum is taken over all finite Borel partitions  of $X$ with diameter at most $\epsilon$.
\item \textbf{Katok's  $\epsilon$-entropy}
\end{itemize}

Let $\mu \in  \mathcal{M}(X)$, $\epsilon>0$, $t>0$ and  $\delta \in (0,1)$.
Put
$$R_\mu^\delta(\phi, t, \epsilon)=\min\{\#E: E\subset X ~\text{and} ~\mu (\cup_{x\in E}B_t(x,\epsilon,\phi))> \delta\}.$$
We define \emph{the upper and lower  Katok' $\epsilon$ entropies of $\mu$ with respect to $\phi$} as 
\begin{align*}
\overline{h}_{\mu}^K(\phi,\epsilon, \delta)&=\limsup_{t\to \infty} \frac{1}{t} \log R_\mu^\delta(\phi, t, \epsilon),\\
\underline{h}_{\mu}^K(\phi,\epsilon, \delta)&=\liminf_{t\to \infty} \frac{1}{t} \log R_\mu^\delta(\phi, t, \epsilon). 
\end{align*}
For time one map, one can  similarly define  the quantities $R_\mu^\delta(\phi_1, n, \epsilon)$,
$\overline{h}_{\mu}^K(\phi_1,\epsilon, \delta), 
\underline{h}_{\mu}^K(\phi_1,\epsilon, \delta)$.

Finally, we give the proof of Theorem  \ref{thm 1.3}.

\begin{proof}[Proof of Theorem \ref{thm 1.3}]
It suffices to show 
\begin{align*}
\over&=\limsup_{\epsilon \to 0}\frac{1}{\logf}\sup_{\mu \in \mathcal{E}_{\phi}(X)}F(\mu,\epsilon).
\end{align*}
For uniformly  Lipschitz flow,  we have  $\over=\ov$ by  Proposition \ref{prop2.4}. We  divide the proof  into  three  steps.

\text{\textbf{Step 1}} We show 
\begin{align*}
\over&=\limsup_{\epsilon \to 0}\frac{1}{\logf}\sup_{\mu \in \mathcal{M}_{\phi}(X)}\inf_{\diam (P) \leq \epsilon}\limits h_\mu(\phi_1,P),\\
&=\limsup_{\epsilon \to 0}\frac{1}{\logf}\sup_{\mu \in \mathcal{E}_{\phi}(X)}\inf_{\diam (P) \leq \epsilon}\limits h_\mu(\phi_1,P).
\end{align*}

One has
\begin{align*}
\over&=\ov\\
&=\limsup_{\epsilon \to 0}\frac{1}{\logf}\sup_{\mu \in \mathcal{M}_{\phi_1}(X)}\inf_{\diam (P) \leq \epsilon}h_\mu(\phi_1,P), \\
&\geq \limsup_{\epsilon \to 0}\frac{1}{\logf}\sup_{\mu \in \mathcal{M}_{\phi}(X)}\inf_{\diam (P) \leq \epsilon}h_\mu(\phi_1,P),\\
& \geq\limsup_{\epsilon \to 0}\frac{1}{\logf}\sup_{\mu \in \mathcal{E}_{\phi}(X)}\inf_{\diam (P) \leq \epsilon}h_\mu(\phi_1,P),
\end{align*}
where the second $``="$ holds by \cite[Theorem 3.1]{gs20}.

Let $\ergodic$. For every $t\in \mathbb{R}$,  we define
\begin{align*}
\mu_t(B):=\mu(\phi_t(B)),
m(B):=\int_{0}^{1} \mu_t(B)dt,
\end{align*}
where  $B$ is a Borel measurable set. Then we have  $\mu_t\in \mathcal{E}_{\phi_1}(X)$ and $m\in\mathcal{E}_{\phi}(X) $. Since  $\phi_{-t}$ is an  invertible measure-preserving transformation  between  $(X,\mu_t,\phi_1)$ and $(X,\mu,\phi_1)$, then 
$h_\mu(\phi_1,\phi_t(\xi))=h_{\mu_t}(\phi_1,\xi)$ holds for any  finite measurable  partition $\xi$. Choose   a  finite  Borel partition $\xi$  of $X$ with diameter  less than $\epsilon/L$.  Recall that $L=L(1)$.  Notice that $-x\log x$ is a  concave function.  By Jensen's inequality,  one has 
$$h_m(\phi_1,\xi)\geq \int_{0}^{1}h_{\mu_t}(\phi_1,\xi)dt.$$
So there is a $t_0\in [0,1]$ so that
$$h_m(\phi_1,\xi) \geq h_{\mu_{t_0}}(\phi_1,\xi)=h_\mu(\phi_1,\phi_{t_0}\xi)\geq \inf_{\diam (P) \leq \epsilon}h_\mu(\phi_1,P ).$$
It follows that
$$ \sup_{\mu\in \mathcal{E}_{\phi_1}(X)}\inf_{\diam (P) \leq \epsilon}\limits h_\mu(\phi_1,P)\leq \sup_{m\in \mathcal{E}_{\phi}(X)}\limits \inf_{\diam \xi \leq \epsilon/L}\limits h_m(\phi_1,\xi).$$
Together with \cite[Remark 3.6]{gs20},  one  has  $$\over\leq \limsup_{\epsilon \to 0}\frac{1}{\logf}\sup_{\mu \in \mathcal{E}_{\phi}(X)}\inf_{\diam (P) \leq \epsilon}h_\mu(\phi_1,P).$$ 
\text{\textbf{Step 2}} We show  that for every $\delta \in (0,1)$, 
\begin{align*}
\over&=\limsup_{\epsilon \to 0}\frac{1}{\logf}\sup_{\mu \in \mathcal{M}_{\phi}(X)}\overline{h}_{\mu}^K(\phi_1,\epsilon, \delta),\\
&=\limsup_{\epsilon \to 0}\frac{1}{\logf}\sup_{\mu \in \mathcal{E}_{\phi}(X)}\overline{h}_{\mu}^K(\phi_1,\epsilon, \delta).
\end{align*} 

We  only prove the equalities for $\overline{h}_{\mu}^K(\phi_1,\epsilon, \delta)$   since similar method is valid for $\underline{h}_{\mu}^K(\phi_1,\epsilon, \delta)$. Fix $\delta \in (0,1)$ and let  $\epsilon >0$.   Using the  fact  that   $R_\mu ^\delta (\phi_1,n,\epsilon) \leq r_n(\phi_1,X,d,\epsilon)$  for every  $\i$  and $n \in\mathbb{N}$, one has 
\begin{align*}
\overline{\rm mdim}_M (\phi, X, d )
&\geq  \limsup_{\epsilon\to 0} \frac{1}{\log \frac{1}{\epsilon}} \sup_{\mu\in \mathcal{M}_{\phi}(X)} \overline{h}_{\mu}^K(\phi_1,\epsilon, \delta)\\
&\geq \limsup_{\epsilon\to 0} \frac{1}{\log \frac{1}{\epsilon}} \sup_{\mu\in \mathcal{E}_{\phi}(X)} \overline{h}_{\mu}^K(\phi_1,\epsilon, \delta).
\end{align*}

Again, let $\ergodic$. For every $t\in \mathbb{R}$,  we define
\begin{align*}
\mu_t(B):=\mu(\phi_t(B)),
m(B):=\int_{0}^{1} \mu_t(B)dt,
\end{align*}
where  $B$ is a Borel measurable set. Then  we have $\mu_t\in \mathcal{E}_{\phi_1}(X)$ and $m\in\mathcal{E}_{\phi}(X) $.  Let $E_n$ be a subset of $X$  so that  $\#E_n= R_{m}^{\delta}(\phi,n+1,\epsilon)$. Then 
$$ m(\cup_{x\in E_n} B_{n+1}(x,\epsilon,\phi))=\int_0^1 \mu_t(\cup_{x\in E_n}B_{n+1}(x,\epsilon,\phi))dt >\delta.$$
Choose  a $t_0\in [0,1]$ so that 
$$\mu_{t_0}(\cup_{x\in E_n}B_{n+1}(x,\epsilon,\phi))=\mu(\cup_{x\in E_n} \phi_{t_0}B_{n+1}(x,\epsilon,\phi))>\delta.$$
Since  $\phi_{t_0}\left(B_{n+1}(x,\epsilon,\phi)\right)\subset B_n(\phi_{t_0}x,\epsilon,\phi_1)$, then 
$\mu(\cup_{x\in E_n}B_n(\phi_{t_0}x,\epsilon,\phi_1))>\delta.$  So  $R_{\mu}^{\delta}(\phi_1,n,\epsilon)\leq R_{m}^{\delta}(\phi, n+1,\epsilon)$ for every $n\in \mathbb{N}$. This  yields that 
\begin{align*}
\overline{h}_{\mu}^K(\phi_1,\epsilon,\delta)\leq
 \overline{h}_{m}^K(\phi,\epsilon,\delta)&\leq \sup_{m\in \mathcal{E}_{\phi}(X)}\overline{h}_{m}^K(\phi,\epsilon,\delta)
 \leq \sup_{m\in \mathcal{E}_{\phi}(X)}\overline{h}_{m}^K(\phi_1,\epsilon/L,\delta).
\end{align*}
By \cite[Proposition 6.3]{shi}, we obtain 
$$\overline{\rm mdim}_M (\phi, X, d )\leq  \limsup_{\epsilon\to 0} \frac{1}{\log \frac{1}{\epsilon}} \sup_{\mu\in \mathcal{E}_{\phi}(X)} \overline{h}_{\mu}^K(\phi_1,\epsilon, \delta).$$

\text{\textbf{Step 3}} We show
\begin{align*}
\over&=\limsup_{\epsilon \to 0}\frac{1}{\logf}\sup_{\mu \in \mathcal{M}_{\phi}(X)}\overline{h}_\mu^{BK}(\phi_1,\epsilon),\\
&=\limsup_{\epsilon \to 0}\frac{1}{\logf}\sup_{\mu \in \mathcal{E}_{\phi}(X)}\overline{h}_\mu^{BK}(\phi_1,\epsilon).
\end{align*} 
  By Theorem \ref{thm 1.2}, we have 
\begin{align*}
\over
&\geq\limsup_{\epsilon \to 0}\frac{1}{\logf}\sup_{\mu \in \mathcal{M}_{\phi}(X)} \overline{h}_\mu^{BK}(\phi,\epsilon),\\
&\geq\limsup_{\epsilon \to 0}\frac{1}{\logf}\sup_{\mu \in \mathcal{M}_{\phi}(X)} \overline{h}_\mu^{BK}(\phi_1,\epsilon),\\
&\geq\limsup_{\epsilon \to 0}\frac{1}{\logf}\sup_{\mu \in \mathcal{E}_{\phi}(X)} \overline{h}_\mu^{BK}(\phi_1,\epsilon).
\end{align*}

Fix $\epsilon>0$ and $\e$. Notice that for any  $s,t\geq0$ and $x\in X$, one has $\phi_s(B_{t+s}(x,\epsilon,\phi)) \subset  B_{t}(\phi_sx,\epsilon,\phi)$. Then the ergodicity of $\mu$ implies that  $\overline{h}_\mu^{BK}(\phi,x,\epsilon)$ is a constant for $\mu$-a.e.$x\in X$,  and this constant is exactly equals to $\overline{h}_\mu^{BK}(\phi,\epsilon)$.   Let  $\delta >0$ and $s>\overline{h}_\mu^{BK}(\phi,\epsilon)$. Put 
$$E_n:=\{x\in X:  \mu(B_k(x,\epsilon,\phi))>e^{-ks}, \forall  k\geq n\}.$$
Then $\mu(\cup_{n\geq 1} E_n)=1$.  By the  continuity of $\mu$, choose $n_0$ so  that  $\mu(E_{n_0})>\delta$ for any $n\geq n_0$ . Let   $n\geq n_0$ and  $F_n \subset E_n$ be a   $(n, 2\epsilon)$-separated set of $E_n$ with  the largest cardinality. Then the open  ball family  $\{B_n(x,\epsilon,\phi), x\in F_n\}$  are pairwise disjoint. This yields that 
$$1\geq \mu(\cup_{x\in F_n}B_n(x,\epsilon,\phi))=\sum_{x\in F_n}\mu (B_n(x,\epsilon,\phi))\geq \# F_n\cdot e^{-ns}.$$
So  $R_\mu ^\delta (\phi,n,2\epsilon)\leq e^{ns}$  for any $n\geq N_0$. Consequently, 
$$\overline h_\mu^K (\phi_1,2\epsilon,\delta)\leq \overline h_\mu^K (\phi, 2\epsilon,\delta)\leq  \overline{h}_\mu^{BK}(\phi,\epsilon)\leq  \overline{h}_\mu^{BK}(\phi_1,\epsilon/L).$$ Using Step 2,  we get 
$$\overline{\rm mdim}_M (\phi, X, d )
\leq \limsup_{\epsilon\to 0} \frac{1}{\log \frac{1}{\epsilon}} \sup_{\mu\in \mathcal{E}_{\phi}(X)}  \overline{h}_\mu^{BK}(\phi_1,\epsilon).$$
\end{proof}

Finally, we finish this paper with three questions as follows.

\text{\bf Question 1}~~Is it possible that the  $"\leq"$ can be strict in Proposition  \ref{prop2.4}?   

\text{\bf Question 2}~~ Does there exist the  metric $d$ compatible with the topology of  $X$ such that $\text{mdim}(X,\phi)= \over$?

\text{\textbf{Question 3}}  Can the condition of uniformly Lipschitz flow  be removed in Theorem \ref{thm 1.3}?

\section*{Acknowledgement} 

\noindent The work was supported by the
National Natural Science Foundation of China (Nos.12071222 and 11971236). The work was also funded by the Priority Academic Program Development of Jiangsu Higher Education Institutions.  We would like to express our gratitude to Tianyuan Mathematical Center in Southwest China(11826102), Sichuan University and Southwest Jiaotong University for their support and hospitality.


\end{document}